# Calibrating Verbal Uncertainty as a Linear Feature to Reduce Hallucinations

### v1

@@ 4.2 Hallucination Mitigation via Inference-time Mechanistic Uncertainty Calibration
  • Confident Hallucination Rate: The proportion of responses that are not entailed by the golden answer and have a low VU below a predefined threshold. The threshold is identified by minimizing the sum of squared distances from VU to the threshold (Kossen et al., 2024).
@@
  • Correct Rate: The proportion of samples aligned with the golden answer.
@@
Hallu.
Rate $\downarrow$
@@
Rate $\downarrow$
Corr.$\uparrow$
@@
Hallu.
Rate $\downarrow$
@@
Rate $\downarrow$
Corr.$\uparrow$
@@
Table 5: Ablation Study Results for Llama-3.1-8B-Instruct when the VUF is extracted from TriviaQA only and it is used to intervene on NQ-Open and PopQA samples.
@@
We compare results before and after the application of MUC with calculated uncertainties on TriviaQA, NQ-Open, and PopQA in Tab. 3. Overall, the application of MUC results in a substantial reduction of confident hallucinations at the cost of a small decrease in Correct Rate. [^7] Different trade offs between Hallucination Rate and Correct Rate can be obtained by varying the strength of the inference-time interventions. The optimal trade-off is highly application-dependent. The observed decrease in the VU/SU Disagreement Rate, along with an increase in the Correlation Coefficient, demonstrates improved calibration and alignment between verbal and semantic uncertainties across most models and datasets. While the VU for incorrect answers increased significantly, indicating reduced confidence in incorrect outputs, the VU for correct answers remained relatively unchanged after uncertainty calibration. [^8] The increase in VU also for correct answers is not necessarily wrong, as it stems from the cases where the model is right, but semantically uncertain. We observe a similar phenomenon in § 3.2, where Qwen2.5 is less sensitive to intervention.

### v2

@@ 4.2 Hallucination Mitigation via Inference-time Mechanistic Uncertainty Calibration
  • Overall Hallucination Rate: The proportion of samples where the model doesn’t refuse and the answer is not entailed by the golden answer.
  • Confident Hallucination Rate: The proportion of responses that are not entailed by the golden answer and have a low VU below a predefined threshold. The threshold is identified by minimizing the sum of squared distances from VU to the threshold (Kossen et al., 2024) [^10] In the original output, confident hallucinations (incorrect and certain) and hallucinations (incorrect and non-refusal) are the same, as verbally uncertain answers without refusal are rare.
  • Correctness Rate: The proportion of samples entailed by the golden answer.
@@
Conf. Hallu.
Rate ↓
@@
Rate ↓
Corr.↑
@@
Conf. Hallu.
Rate ↓
@@
Rate ↓
Corr.↑
@@
w/ VUF 21.6 44.7 0.34 0.73 0.20
Table 5: Ablation Study Results for Llama-3.1-8B-Instruct when the VUF from TriviaQA, NQ-Open, and PopQA datasets is replaced with: (1) VUF extracted from TriviaQA only, applied to intervene on NQ-Open and PopQA samples. (2) random vectors, applied to intervene on three datasets.
@@
We compare results before and after the application of MUC with calculated uncertainties on TriviaQA, NQ-Open, and PopQA in Tab. 3. Overall, the application of MUC results in a substantial reduction of confident hallucinations at the cost of a small decrease in Correctness Rate.  [^11] The correctness rate decreases when the correct answer exhibits high SU and the MUC intervention results in an abstention. This reduction stems from noise in measuring SU. For the same question, the sampled answers may contain diverse additional information, leading to a high calculated semantic entropy. Please see Appendix 13.3 for an example. Different trade offs between Hallucination Rate and Correct Rate can be obtained by varying the strength of the inference-time interventions. The optimal trade-off is highly application-dependent. The observed decrease in the VU/SU Disagreement Rate, along with an increase in the Correlation Coefficient, demonstrates improved calibration and alignment between verbal and semantic uncertainties across most models and datasets. While the VU for incorrect answers increased significantly, indicating reduced confidence in incorrect outputs, the VU for correct answers remained relatively unchanged after uncertainty calibration. [^12] The increase in VU also for correct answers is not necessarily wrong, as it stems from the cases where the answer is correct, but the model is semantically uncertain about it. The fact that the trend is consistent across different models and sizes underscores the generality and effectiveness of our approach.
@@
To further prove the importance of using VUF in the MUC method, we perturb activations with random vectors with the same $\alpha$ and value range as the VUF, when the hallucination detector trigger. As shown in Tab. 5, applying random perturbations upon hallucination detection does lead to an improvement over the baseline. However, the results show that intervening on the VUF direction is significantly more effective than random perturbations.